OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 3x – 5y + z = 12 x + 6y – z = –4 2x – 5y – 2z = –8 4x + y = 8 –17y = 0 2x – 11y = 16 4x + 7y = –16 –23y + 4z = 24 –17y = 0 13y – 2z = 24 7y – 4z = 0

OpenStudy (anonymous):

You have to get y by itself, i would recommend

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

add 5y to both sides, getting 3x+z=12+5y. I usually just find the answer to the first equation and then see if it works in any of the others. From 3x+z=12+5y, you subtract 12 from both sides to get 3x+z-12=5y. Then, you divide both sides by 5, getting 3/5x+1/5z-12/5=y. Solve from there.

OpenStudy (anonymous):

I know that the second system in x term would be x=-4-6y+7

OpenStudy (anonymous):

well these are the answer choices 4x + y = 8 –17y = 0 2x – 11y = 16 4x + 7y = –16 –23y + 4z = 24 –17y = 0 13y – 2z = 24 7y – 4z = 0

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

what are the top three equations

OpenStudy (anonymous):

3x – 5y + z = 12 x + 6y – z = –4 2x – 5y – 2z = –8

OpenStudy (anonymous):

i think I get it just give me a moment

OpenStudy (anonymous):

my guess and this is a guess because i am not getting the exact answer would be the -23y+4z=24 and -17y=0